40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 28 04:57:00 2011, kiwamu, Update, LSC, POX11 demod board broken Screen_shot_2011-10-28_at_2.26.02.png
    Reply  Thu Nov 3 18:41:36 2011, kiwamu, Update, LSC, POX11 demod board : haven't been modified yet 
       Reply  Thu Nov 3 19:58:18 2011, kiwamu, Update, LSC, POX11 demod board : modification done POX11demod.png
          Reply  Fri Nov 4 00:13:24 2011, Koji, Update, LSC, XARM locked with POX11 POX11.pdf
Message ID: 5753     Entry time: Fri Oct 28 04:57:00 2011     Reply to this: 5801
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: POX11 demod board broken 
The POX11 demodulation board is broken. It needs to be fixed in the daytime tomorrow.
It only outputs the Q signal and nothing is coming out from the I output.


(Some stuff checked)
 [OK] ADCs
 [OK] Whitening filters looked fine. Their gains were controllable from EPICS.
 [OK] Connection between the POX11 RFPD and demodulator box
 [OK] The Q signal showed the PDH signal of the X arm with an amplitude of about 200 counts, which almost is the same as that of POY11.
 [NOT GOOD] The I signal from the demod board
 [OK] Outputs cables, which send the I and Q signals from the demod board to WFs are fine.
         
(Test on the POX11 demod board)
 As usual, a test signal whose frequency is shifted by a little bit from that of LO was injected to the RF input of the board to see if the circuit is working.
The I signal didn't show up and there were no signals even in the monitor LEMO output.
Something is wrong in the I signal demodulation path on the circuit board.

Here is an actual time series of the I and Q signals in dataviewer. The I signal outputs just junk while the Q showed a nice sine curve.
Screen_shot_2011-10-28_at_2.26.02.png

ELOG V3.1.3-